OpenStudy (anonymous):

what is the length magnitude of the vector 2,-1?

myininaya (myininaya):

\[|v|=|<a.b>|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

what is the the formula, and what a and b i dont understand what you just wrote

myininaya (myininaya):

a=2 and b=-1 where v is the vector <a,b>=<2,-1>

myininaya (myininaya):

just plug in 2 for a and -1 for b

myininaya (myininaya):

and you will be done :) after simplifying

OpenStudy (anonymous):

1.73?

myininaya (myininaya):

\[\sqrt{(2)^2+(-1)^2}=\sqrt{4+1}=\sqrt{5}\]

OpenStudy (anonymous):

sqrt 5

OpenStudy (anonymous):

a vector has both a magnitude and a ? is it a number?

myininaya (myininaya):

direction